OpenStudy (hoothoot23):

How many solutions does the equation 3x − 7 = 4 + 6 + 4x have?

OpenStudy (hoothoot23):

@jim_thompson5910

OpenStudy (dmezzullo):

How many do you think it has?

OpenStudy (hoothoot23):

infinitely many

OpenStudy (dmezzullo):

Do you have answer choices for this?

OpenStudy (hoothoot23):

yes

OpenStudy (hoothoot23):

How many solutions does the equation 3x − 7 = 4 + 6 + 4x have? Two None Infinitely many One

OpenStudy (mecharv):

Is it specified whether it should be integral or something?

jhonyy9 (jhonyy9):

3x-7 = 4+6+4x -7-4-6=4x-3x -17 = x

OpenStudy (hoothoot23):

so what is in one

OpenStudy (hoothoot23):

i mean is it one

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

yes one solution x=-17

OpenStudy (mathmale):

"3x − 7 = 4 + 6 + 4x " being a first-order linear equation, we'd expect ONE root. Consolidate the 2 x-terms on the right side of this equation and the constant terms on the left side. Solve for x. check your answer via substitution.